OpenStudy (anonymous):

using an identity, (sinx+cosx)^2=1+sin2x

OpenStudy (anonymous):

(sinx+cosx)^2=1+sin2x "=1+2sinxcosx "=sin^2x+cos^2x+2sinxcosx "=(sinx+cosx)^2